OpenStudy (anonymous):

claire has received scores of 85, 88, 87, and 85 on her tests. to have an overall score of at least 88 what must she score on her fifth test?

OpenStudy (amistre64):

\[\frac{old.sum+new}{(4+1)tests}=88\] \[old.sum+new=88(5.tests)\] \[new=88(5)-old.sum\]

OpenStudy (anonymous):

so i'm getting 86.6...?

OpenStudy (amistre64):

new = 88(5)-345 so i doubt its a decimal

OpenStudy (amistre64):

id say 95

OpenStudy (anonymous):

aaah, got it! yep 95!